OpenStudy (anonymous):

Write an equation in point-slope form that fits the description given. 1) contains the point (5,8) and is parallel to y = 5x + 12 2) contains the point (6,1) and is perpendicular to y = 3x - 22 3) contains the point (2,5) and is parallel to y = 8x + 10

OpenStudy (anonymous):

What is the correlation b/w parallel lines and the slope? b/w perpendicular lines and the slope?

OpenStudy (anonymous):

For each one: 1. Find the slope of the given line 2. based on what type of line you are looking for - what is the slope of the new line 3. use point slope form to find the new line y - y1 = m(x - x1) plug in your point of x1,y1 and new slope for m rewrite this into y = mx + b format

OpenStudy (anonymous):

you catch all that?

OpenStudy (anonymous):

It's a bit confusing for me

OpenStudy (anonymous):

i didn't think so 1) contains the point (5,8) and is parallel to y = 5x + 12

OpenStudy (anonymous):

the " parallel to y = 5x + 12 " says the slope is the same as the slope of \(y=5x+12\) which looks just like \(y=mx+b\) a line with slope \(m\) so your slope is \(5\)

OpenStudy (anonymous):

I didn't think I was too convoluted.

OpenStudy (anonymous):

Do you understand the correlation b.w slope and types of lines...that is very important, hence my first question. But you are in good hands! Happy studying

OpenStudy (anonymous):

the slope is 5, the point is \((5,8)\) write it as \[y-y_1=m(x-x_1)\]with \[m=5, x_1=5,y_1=8\] you get \[y-8=5(x-5)\]

OpenStudy (anonymous):

I'm not sure about the correlation but I'll try to figure it out thanks

OpenStudy (anonymous):

@JuanitaM it was not at all convoluted, but the steps may not be clear

OpenStudy (anonymous):

Yeah, It didn't sound clear to me thats why

OpenStudy (anonymous):

you good from \[y-8=5(x-5)\] because you are still not done

OpenStudy (anonymous):

Yeah, I got y−8=5(x−5)

OpenStudy (anonymous):

i mean after that?

OpenStudy (anonymous):

steps are always the same multiply out on the right using the distributive law \[y-8=5x-25\] then add \(8\) to get \[y=5x-17\]

OpenStudy (anonymous):

Alright, thanks. Working on it right now

OpenStudy (anonymous):

2) contains the point (6,1) and is perpendicular to y = 3x - 22

OpenStudy (anonymous):

perpendicular to y = 3x - 22 tells you it is perpendicular to a line with slope \(3\) that means the slope of your line is the "negative reciprocal" i.e \[-\frac{1}{3}\]

OpenStudy (nikato):

actually, you can stop at y−8=5(x−5), which is already in point-slope form if you go further,y=5x-17, this is slope-intercept form

OpenStudy (anonymous):

So I can just stop at y - 8 = 5 (x-5) ?

OpenStudy (anonymous):

if it says "point slope" then yes, stop

OpenStudy (anonymous):

similarly \[y-1=-\frac{1}{3}(x-6)\] for the second one